OpenStudy (anonymous):

Find lim_{x rightarrow 0} (tan 3x^{2} + sin ^{2} 5x) / x ^{2} i'd tried applying L'Hospital Rule twice, and i'd obtained -25 as the answer. is there any other method to solve this question? or if there aint any, is my answer correct?

OpenStudy (turingtest):

\[\frac{\tan(3x^2)+\sin^2(5x)}{x^2}\]?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} (\tan 3x^{2} + \sin^{2} 5x) / x ^{2} \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8, I believe

OpenStudy (turingtest):

I get a mess...

OpenStudy (akshay_budhkar):

\[\lim x \rightarrow0 {tanx \over x} = 1 and \lim x \rightarrow 0{sinx \over x} = 1 \] also when x tends to zero x^2 tends to zero

OpenStudy (akshay_budhkar):

so you can use that limit to get the answer as\[\lim x \rightarrow 0 {\tan (3x^2) \over x^2 } = \lim x \rightarrow0(3{\tan(3x^2) \over 3x^2}) = 3\]

OpenStudy (anonymous):

Break it up into parts. First, the limit of \[ \frac{\tan(3x^2)}{x^2} \] As the limit goes to zero, tangent behaves just like sine. The Maclaurin series would therefore begin with 3x^2. The x^2 terms cancel, leaving a 3. Then, \[\frac{\sin^2(5x)}{x^2} = \left(\frac{\sin(5x)}{x}\right)^2 \] the limit of which is 25. So I should revise my answer, I believe it should be 28.

OpenStudy (akshay_budhkar):

o_O?

OpenStudy (turingtest):

wolf agrees with you Jemurray3!

OpenStudy (akshay_budhkar):

umm so where did i go wrong?

OpenStudy (anonymous):

Hooray! @Akshay I don't think you did. The limit of the tangent part is 3, as you say, and the limit of the sine part is 5^2 = 25

OpenStudy (akshay_budhkar):

oh yea! i am still to solve the sine part -_-

OpenStudy (akshay_budhkar):

my bad :P

OpenStudy (anonymous):

@.@ wow means i need to work harder on calculus == anyway i handed in the answer last week. not expecting to obtain any solutions from my lecturer, that's why i asked here. he's just touching the tip of the topic... so conclusion is working out separately to find the limits?

OpenStudy (anonymous):

Also, I should apologize.... that wasn't a very academic answer. Filled with my own little shortcuts and intuitive reasoning. To be more precise, notice that \[ \frac{\tan(3x^2)}{x^2} = 3\frac{\tan(3x^2)}{3x^2} = 3\frac{\tan(u)}{u} \] The limit of that as u -> 0 is 3.

OpenStudy (anonymous):

Just like Akshay said. Then, \[ \frac{sin^2(5x)}{x^2} = \left( 5\frac{\sin(5x)}{5x}\right)^2 = 25\left(\frac{\sin(5x)}{5x}\right)^2\] \[ = 25\left(\frac{\sin(u)}{u} \right)^2 \] The limit of which as u -> 0 is 25.

OpenStudy (anonymous):

owh. thank you all very much. need some time to read up. lolx.

OpenStudy (anonymous):

off topic: im 20yo this year, am i too stupid for not knowing how to solve this question? ==

OpenStudy (akshay_budhkar):

Age is just a number, it cant define what you should know and what you shouldnt

OpenStudy (akshay_budhkar):

Also stupidity is a relative concept, depends on who looks at it in what manner. Even if a 50 year old didn't know this , i would not tell he is stupid. The fact that you know how to do it now is of more importance

OpenStudy (anonymous):

mmm could be my foundation on trigo and calculus aint strong enough. i always hated trigo lol

OpenStudy (akshay_budhkar):

lol the foundation can be built anytime :P